OpenStudy (anonymous):

What is cos(2∂) equal to?

OpenStudy (ash2326):

\[\cos (2\delta)= 2\cos^2 (\delta)-1\]

OpenStudy (vishal_kothari):

\[\cos ^{2}\theta-\sin ^{2}\theta\]

OpenStudy (anonymous):

thnx! so does 2cos2(δ)−1=cos2θ−sin2θ?

OpenStudy (ash2326):

or \(cos^2\delta-sin^2 \delta\) or \(1-2sin^2\delta\)